OpenStudy (horsegal244):

A scale on a map shows that 2 inches = 25 miles. Part A How many inches on the map represents 60 miles? Part B How many miles is represented by 1/4 inch on the map? Medals And Fan!

OpenStudy (horsegal244):

@pooja195

OpenStudy (aaronandyson):

A) 2 inches=25 miles x inches = 60 miles B) 2 inches=25 miles 1/4 inches= x miles

OpenStudy (horsegal244):

i think that part A is 5 inches

OpenStudy (horsegal244):

@AaronAndyson

OpenStudy (aaronandyson):

Yes?

OpenStudy (horsegal244):

I think that part A is 5 inches.

OpenStudy (aaronandyson):

2*60 = 120 120/25 = 4.8

OpenStudy (horsegal244):

Ok so about 5

OpenStudy (aaronandyson):

Yes.